OpenStudy (richyw):

could someone give me a refresher on what to do with somthing like this \[\sin\left( 2\sin^{-1}\left( \frac{1}{2}\right)\right)\]

OpenStudy (anonymous):

Start from inside out: 1/2 = sin 30° -> 2arcsin1/2 = 2 * 30 = ... => sin (.. ) = ...

OpenStudy (richyw):

thanks, I made a typo though. I meant to say \[\sin\left( 2\sin^{-1}\left( \frac{x}{2}\right)\right)\]

OpenStudy (richyw):

\[y=\sin\left( 2\sin^{-1}\left( \frac{x}{2}\right)\right)\]\[\sin^{-1}y=2\sin^{-1}\left(\frac{x}{2}\right)\]

OpenStudy (richyw):

nope. super lost!

OpenStudy (anonymous):

sin2(theta) = sin(theta)*cos(theta)

OpenStudy (anonymous):

Use the half angle formula: sin2a = 2 sina cosa with a = arcsin (x/2)

OpenStudy (anonymous):



OpenStudy (anonymous):

so x/2 * (sqrt(4-x^2))/2

OpenStudy (richyw):

thanks! totally forgot how useful those triangle diagrams are!